OpenStudy (pphalke):

A 14-foot piece of string is cut into two pieces so that the longer piece is 2 feet longer than twice the shorter piece. if the shorter piece is x feet long. find the lengths of both pieces

OpenStudy (faiqraees):

Short peice = x Long piece = 2x+2 Total =14 x+2x+2=14

OpenStudy (pphalke):

is the answer to the last one 2x+2=14?

OpenStudy (pphalke):

faiqraees

OpenStudy (pphalke):

how else do I solve after this

OpenStudy (yumyum247):